# LLM-SmartAudit: Advanced Smart Contract Vulnerability Detection

### v1

@@ III-B2 Execution Mode
The LLM-SmartAudit task queue comprises three primary subtasks: Contract Analysis, Vulnerabilities Identification, and Security Summary. Each subtask employs task-oriented role-playing, involving two distinct roles collaborating to achieve specific objectives. Initially, the Project Manager establishes the primary goal for the team and collaborate with the Smart Contract Auditor to assess the contract’s purpose and structure. The Smart Contract Counselor then reviews and summarizes these initial findings. This preliminary analysis, along with the smart contract codes, is then forwarded to the next phase. In the next subtask, the Smart Contract Auditor and the Solidity Programming Expert work together to identify security weaknesses within the contract. In the final subtask, the Smart Contract Auditor and the Solidity Programming Expert jointly formulate a comprehensive analysis report, detailing all identified vulnerabilities and their potential impacts.
Figure 3: Task Queue in BA mode and TA mode
@@
To evaluate the effectiveness of this approach, LLM-SmartAudit implements two distinct operational strategies, as illustrated in Figure 3. These strategies, namely Broad Analysis (BA) and Targeted Analysis (TA).
@@
BA mode, built on zero-shot-CoT prompting, emphasizes a comprehensive examination of smart contracts. It utilizes the combined expertise of the Smart Contract Auditor and the Solidity Programming Expert, with strategic guidance from the Counselor. This mode leverages the LLMs’ capabilities to identify a wide spectrum of potential vulnerabilities. BA mode is particularly effective for its versatility and adaptability in handling diverse vulnerability detection tasks. To prevent potential deadlocks where agents fail to reach a consensus on vulnerability assessments, LLM-SmartAudit limit the maximum number of interaction rounds between them to a predefined number $n$.
@@
TA mode, built on the few-shot-CoT prompting, adopts a scenario-based approach to vulnerability detection, dividing the Vulnerabilities Identification phase into 40 distinct scenarios, each targeting a specific vulnerability type vulnerability types are listed in our resposity. This mode leverages collaborative efforts between the Smart Contract Auditor and the Solidity Programming Expert, who guide the model in detection techniques and provide relevant examples. This structured approach ensures targeted analysis and seamless information flow, effectively utilizing examples to direct the model towards identifying specific vulnerabilities.
III-C Collaborative Decision-Making
@@ III-C1 Collaborative Analysis
(a) Conversation-Driven Control Flow
(b) A Simplified Depiction of the Analysis Procedure
Figure 4: Collaborative Decision-Making between Two Agents
(a) Simple Question-Answering
(b) Role Exchanges Question-Answering
Figure 5: Examples for Different Question-Answering
Figure 4(a) illustrates the conversation-driven control flow, demonstrating how task processes are executed between two agents through conversations. The process begins with the user agent sending a prompt to the assistant agent, which then generates a response via the language model API. The response is relayed back to the user agent, which generates a new prompt within the multi-conversation system. This new prompt is then sent to the assistant agent, initiating the next round of analysis.
Figure 4(b) illustrates a simplified example of the smart contract analysis procedure. In this example, the system analyzes a contract potentially affected by an Arithmetic vulnerability. The process begins with the Project Manager initiating the audit task, setting the team’s objective and initiating the discussion about the contract under review. As the analysis progresses, the Smart Contract Auditor and Project Manager contribute their expertise, sharing insights on the contract’s purpose and structure. Finally, the Smart Contract Counselor summarizes the initial findings and provides a phase report. In the next phase, the Solidity Programming Expert provides a detailed code analysis, which the Auditor uses to identify potential Integer overflow/underflow vulnerabilities, offering a comprehensive description. The process concludes with the Counselor compiling a comprehensive audit report, summarizing all identified vulnerabilities and their potential impacts.
III-C2 Role Exchanges
Traditional LLM can sometimes produce inaccuracies or irrelevant information, especially in complex tasks like generating code insights. For example, as illustrated in Figure 5(a), if the Smart Contract Auditor is instructed to review a vulnerable contract code previously identified with Arithmetic vulnerabilities, there’s a risk of receiving misleading feedback. In such cases, the Auditor might erroneously flag the contract as vulnerable to both "Integer Overflow/Underflow" and "Reentrancy", incorrectly generating false positives for Reentrancy. To address this issue, LLM-SmartAudit introduces a roles-swaping mechanism to enhance precision in vulnerability detection.
@@ IV-A Research Questions
TABLE I: Evaluation of Smart Contract Vulnerability Detection Tools — A Comparative Analysis
@@ IV-B Experimental Settings
IV-B1 Benchmarking Dataset
To comprehensively measure the system’s capabilities, robust evaluation datasets are essential. Zhang et al. [37] categorize smart contract vulnerabilities into ‘machine-auditable’ and ‘machine-unauditable’ types. Traditional vulnerability detection tools can detect machine-auditable vulnerabilities, whereas machine-unauditable vulnerabilities require expert human intervention. In this study, machine-auditable vulnerabilities are termed ‘specific vulnerabilities’, while machine-unauditable vulnerabilities are designated as ‘complex logic vulnerabilities’.
@@ IV-B1 Benchmarking Dataset
Real-world dataset comprises both specific and complex logic vulnerabilities that have led to actual exploits. This dataset is derived from reputable Code4rena contests [31, 37], which attract global experts and companies to identify vulnerabilities in real-world smart contract projects. Participants receive financial compensation for their discoveries, lending credibility to the reports and confirming that the identified vulnerabilities genuinely reflect real-world attack scenarios. The Real-world dataset contains 102 projects and 6,454 contracts, encompassing 499 high-risk, 909 medium-risk, 1,420 low-risk, and 2,417 ground-level vulnerabilities. The severity classification of these vulnerabilities is based on their potential financial impact on the contract. This study primarily focuses on high-risk and medium-risk vulnerabilities.
IV-B2 Evaluation Criteria
@@ IV-B3 Hardware Configuration
TABLE II: Comparative Evaluation of Smart Contract Vulnerability Detection Across GPT-3.5 and GPT-4 Models
@@ IV-C Experimental Results
IV-C1 RQ1: Comparative Evaluation with Leading Traditional Tools
To address RQ1, we evaluated the effectiveness of LLM-SmartAudit in its BA mode, utilizing the GPT-3.5 model. The study evaluates the tool’s capability to identify 10 specific vulnerabilities types in smart contracts, comparing its performance against leading traditional tools in the domain. The comparative analysis included ten prominent smart contract analysis tools across various categories: formal verification (Securify [38], VeriSmart [39]), symbolic execution (Mythirl [40], Oyente [1]), fuzzing (ConFuzzius [41], sFuzz [42]), intermediate representation (IR) analysis (Slither [43], Conkas [44]), and machine learning approaches (e.g., GNNSCVD [45], Eth2Vec [46]). Table I presents detailed performance metrics for each analyzed tool, specifically focusing on False Positives (FPs) and recall rates.
@@ IV-C1 RQ1: Comparative Evaluation with Leading Traditional Tools
IV-C2 RQ2: Performance of LLM-SmartAudit Across Different Strategies
The analysis conducted for RQ1 revealed specific limitations of LLM-SmartAudit in BA mode, particularly in detecting TOD vulnerabilities. To address these limitations and answer RQ2, LLM-SmartAudit was adapated to the TA mode, and the LLM was upgraded to the more advanced GPT-4 model. Additionally, we introduced a baseline comparison using a zero-shot prompt approach. Table II presents the comprehensive performance metrics for each strategy, including TP, FP, FN, TN, and F1-score.
@@ IV-C2 RQ2: Performance of LLM-SmartAudit Across Different Strategies
TABLE III: Performance Comparison of Different Detection Methods on Real-World Dataset
@@
IV-C3 RQ3: Evaluation Using Real-world Dataset
To address RQ3, we present a comprehensive evaluation of LLM-SmartAudit system using the Real-world dataset. The LLM-SmartAudit system design allocates approximately 1,000 tokens for prompt engineering, excluding the context of code snippets. Given GPT-3.5’s default token limit of 4,096, only 3,000 tokens remain available for contract content within the prompt. For the Labeled dataset, the analyzed contracts are less than 3,000 tokens in length. However, with the Real-world datasets, 533 contracts exceed 3,000 tokens in length. Consequently, only GPT-4 was utilized for this evaluation due to its higher token support of up to 128,000 tokens in context. The evaluation focuses on two primary metrics: TPs and recall, with results compared against audit reports [^3] https://github.com/ZhangZhuoSJTU/Web3Bugs/tree/main/reports. Table III presents the evaluation results.
@@ IV-C3 RQ3: Evaluation Using Real-world Dataset
While earlier works such as David et al. [47] and GPTScan [48] demonstrated the potential of GPT-based vulnerability detection tools, our research significantly broadens the scope and improves upon their findings. However, due to the unavailability of aforementioned tools and insufficient information for replication, this study relies on their published statistics for comparison. David et al. reported a recall rate of 39.73%, detecting 58 out of 146 vulnerabilities using a combination of Claude and GPT. GPTScan analyzed 232 vulnerabilities across 72 projects, correctly identifying 40 true positives. In contrast, our study substantially broadens the scope, examining 1,408 vulnerabilities across 102 projects. This expanded scope allows for a more comprehensive assessment of GPT-based tools’ capabilities in real-world scenarios. Our results show a notable improvement, with 672 out of 1,408 vulnerabilities detected, yielding an overall recall rate of 47.73%.
@@
IV-C4 RQ4: Newly Discovered Vulnerabilities
In response to RQ4, our methods successfully identified 11 vulnerabilities across 4 different types that were not detected in the audit reports from real-world datasets. These findings have been submitted to the Code4rena community for verification. Table IV provides the list of these newly discovered vulnerabilities.
TABLE IV: Summary of Potential Bugs
Issue Name Location
Unlimited Token Approval function ‘supplyTokenTo’ in SushiYieldSource.sol, the line ‘.approve(_sushiRouter, type(uint256).max)’ in NFTXStakingZap.sol, the line ‘_par.approve(address(_a.parallel().core()), uint256(-1))’ in PARMinerV2.sol
Insufficient Input Validation function ‘setTransferRatio’ in sYETIToken.sol
Handling Partial Withdrawals function ‘withdraw’ in yVault.sol, function ‘processWithdraw’ in synthVault.sol
Unchecked External Calls function ‘earn’ and ‘withdraw’ in yVault.sol, line ‘vault.addValue(_amount, msg.sender, address(this));’ and function ‘getUnifiedAssets’ in IndexTemplate.sol, function ‘collectEarnings’, ‘_push’, and ‘_pullUniV3Nft’ in UniV3Vault.sol
@@ IV-D Related Work
David et al. [47] examined the efficacy of LLMs, like GPT-4 and Claude, in auditing DeFi smart contract security. Their study employed a binary classification approach, asking the LLMs to determine whether a contract is vulnerable. Although GPT-4 and Claude demonstrated high true positive rates, they also exhibited significant false positive rates. The researchers highlighted the substantial evaluation cost, approximately 2,000 USD, for analyzing 52 DeFi attacks. Chen et al. [49] performed a comparative analysis of GPT’s smart contract vulnerability detection capabilities against established tools. Their results revealed varying GPT effectiveness across common vulnerability types, encompassing 8 types compared to the 10 in our study. Sun et al. [48] evaluated GPT’s function vulnerability matching using a binary response format (’Yes’ or ’No’) for predefined scenarios. They also highlighted potential false positives due to GPT’s inherent limitations. While their study found no significant improvements with GPT-4, our research and others have demonstrated GPT-4’s enhanced detection capabilities.
In addition to SOTA commercial products, open-source alternatives have been considered for smart contract analysis. Shou et al. [50] integrates Llama-2 model into the process of fuzzing to detect vulnerabilities in smart contracts, aiming to address inefficiencies in traditional fuzzing methods. However, this approach’s efficacy depends on LLMs’ accurate and nuanced understanding of smart contracts, and it confronts challenges in complexity, cost, and dependence on static analysis. Sun et al. [51] explored open-source tools such as Mixtral and CodeLlama against GPT-4 for detecting smart contract vulnerabilities. They discovered that GPT-4, leveraging its advanced Assistants’ functionalities to effectively utilize enhanced knowledge and structured prompts, significantly outperformed Mixtral and CodeLlama. However, the assessment was limited to demonstrations from the Replicate website, potentially not fully representing these LLMs’ capabilities.
V Discussions
V-A Summary of Findings
@@
V-B Threats of Validity
@@
VI Conclusions

### v2

@@ 3.2.2. Execution Mode
The LLM-SmartAudit task queue comprises three primary subtasks: Contract Analysis, Vulnerabilities Identification, and Comprehresive Report. Each subtask employs task-oriented role-playing, involving two distinct roles collaborating to achieve specific objectives. Initially, the Project Manager establishes the primary goal for the team and collaborate with the Smart Contract Auditor to assess the contract’s purpose and structure. The Smart Contract Counselor then reviews and summarizes these initial findings. This preliminary analysis, along with the smart contract codes, is then forwarded to the next phase.
Figure 3. Task Queue in BA mode and TA mode
In the next subtask, the Smart Contract Auditor and the Solidity Programming Expert work together to identify security weaknesses within the contract. In the final subtask, the Smart Contract Auditor and the Solidity Programming Expert jointly formulate a comprehensive analysis report, detailing all identified vulnerabilities and their potential impacts.
@@
To assess the effectiveness of this workflow, LLM-SmartAudit operates in two distinct modes, illustrated in Figure 3: Broad Analysis (BA) and Targeted Analysis (TA).
@@
BA mode, based on the Thought-Reasoning approach, focuses on a thorough, comprehensive examination of smart contracts. In this mode, the expertise of the Smart Contract Auditor and Solidity Programming Expert is augmented with strategic guidance from the Counselor, leveraging the model’s capacity to identify a wide range of potential vulnerabilities. BA mode is particularly effective for its adaptability in handling various types of vulnerability detection tasks. To prevent potential impasses, where agents might struggle to reach a consensus on certain vulnerability assessments, LLM-SmartAudit restricts the maximum number of interaction rounds to a predefined number, $n$.
@@
TA mode, based on the Buffer-Reasoning prompt, employs a scenario-specific approach for vulnerability detection. It divides the Vulnerability Identification phase into 40 targeted scenarios, each focused on a specific vulnerability type (as listed in our repository). This mode harnesses the collaborative efforts of the Smart Contract Auditor and Solidity Programming Expert, who provide the model with detection techniques and relevant examples. The structured, scenario-based framework ensures targeted analysis and smooth information flow, using specific examples to guide the model toward identifying distinct vulnerabilities.
3.3. Collaborative Decision-Making
@@ 3.3.1. Collaborative Analysis
Figure 4. Collaborative Decision-Making between Two Agents
Figure 4 illustrates the conversation-driven control flow and automated agent chat. The conversation-driven control flow demonstrates how task processes are executed between two agents through conversations. The process begins with the user agent sending a prompt to the assistant agent, which then generates a response via the language model API. The response is relayed back to the user agent, which generates a new prompt within the multi-conversation system. This new prompt is then sent to the assistant agent, initiating the next round of analysis.
The automated agent chat illustrates a simplified example of the smart contract analysis procedure. In this example, the system analyzes a contract potentially affected by an Arithmetic vulnerability. The process begins with the Project Manager initiating the audit task, setting the team’s objective and initiating the discussion about the contract under review. As the analysis progresses, the Smart Contract Auditor and Project Manager contribute their expertise, sharing insights on the contract’s purpose and structure. Finally, the Smart Contract Counselor summarizes the initial findings and provides a phase report. In the next phase, the Solidity Programming Expert provides a detailed code analysis, which the Auditor uses to identify potential Integer overflow/underflow vulnerabilities, offering a comprehensive description. The process concludes with the Counselor compiling a comprehensive audit report, summarizing all identified vulnerabilities and their potential impacts.
3.3.2. Role Exchanges
Traditional LLM can sometimes produce inaccuracies or irrelevant information, especially in complex tasks like generating code insights. For example, as illustrated in Figure 5(a), if the Smart Contract Auditor is instructed to review a vulnerable contract code previously identified with Arithmetic vulnerabilities, there’s a risk of receiving misleading feedback. In such cases, the Auditor might erroneously flag the contract as vulnerable to both ”Integer Overflow/Underflow” and ”Reentrancy”, incorrectly generating false positives for Reentrancy. To address this issue, LLM-SmartAudit introduces a roles-swaping mechanism to enhance precision in vulnerability detection.
(a) Simple QA
(b) Role Exchanges QA
Figure 5. Examples for Different Question-Answering
@@ 4.1. Research Questions
Table 1. Evaluation of Smart Contract Vulnerability Detection Tools — A Comparative Analysis
@@ 4.2. Experimental Settings
4.2.1. Benchmarking Dataset
To comprehensively measure the system’s capabilities, robust evaluation datasets are essential. Zhang et al. (Zhang et al., 2023b) categorize smart contract vulnerabilities into ‘machine-auditable’ and ‘machine-unauditable’ types. Traditional vulnerability detection tools can detect machine-auditable vulnerabilities, whereas machine-unauditable vulnerabilities require expert human intervention. In this study, machine-auditable vulnerabilities are termed ‘specific vulnerabilities’, while machine-unauditable vulnerabilities are designated as ‘complex logic vulnerabilities’.
@@ 4.2.1. Benchmarking Dataset
Real-world dataset comprises both specific and complex logic vulnerabilities that have led to actual exploits. This dataset is derived from reputable Code4rena contests (cod, 2023; Zhang et al., 2023b), which attract global experts and companies to identify vulnerabilities in real-world smart contract projects. Participants receive financial compensation for their discoveries, lending credibility to the reports and confirming that the identified vulnerabilities genuinely reflect real-world attack scenarios. The Real-world dataset contains 102 projects and 6,454 contracts, encompassing 499 high-risk, 909 medium-risk, 1,420 low-risk, and 2,417 ground-level vulnerabilities. The severity classification of these vulnerabilities is based on their potential financial impact on the contract. This study primarily focuses on high-risk and medium-risk vulnerabilities.
4.2.2. Evaluation Criteria
@@ 4.2.3. Hardware Configuration
Table 2. Comparative Evaluation of Smart Contract Vulnerability Detection Across GPT-3.5 and GPT-4 Models
@@ 4.3. Experimental Results
4.3.1. RQ1: Comparative Evaluation with Leading Traditional Tools
To address RQ1, we evaluated the effectiveness of LLM-SmartAudit in its BA mode, utilizing the GPT-3.5 model. The study evaluates the tool’s capability to identify 10 specific vulnerabilities types in smart contracts, comparing its performance against leading traditional tools in the domain. The comparative analysis included ten prominent smart contract analysis tools across various categories: formal verification (Securify (Tsankov et al., 2018), VeriSmart (So et al., 2020)), symbolic execution (Mythirl (development team, 2023), Oyente (Atzei et al., 2017)), fuzzing (ConFuzzius (Torres et al., 2021), sFuzz (Nguyen et al., 2020)), intermediate representation (IR) analysis (Slither (Feist et al., 2019), Conkas (Veloso, 2021)), and machine learning approaches (e.g., GNNSCVD (Zhuang et al., 2021), Eth2Vec (Ashizawa et al., 2021)). Table 1 presents detailed performance metrics for each analyzed tool, specifically focusing on False Positives (FPs) and recall rates.
@@ 4.3.1. RQ1: Comparative Evaluation with Leading Traditional Tools
4.3.2. RQ2: Performance of LLM-SmartAudit Across Different Strategies
The analysis conducted for RQ1 revealed specific limitations of LLM-SmartAudit in BA mode, particularly in detecting TOD vulnerabilities. To address these limitations and answer RQ2, LLM-SmartAudit was adapated to the TA mode, and the LLM was upgraded to the more advanced GPT-4 model. Additionally, we introduced a baseline comparison using a zero-shot prompt approach. Table 2 presents the comprehensive performance metrics for each strategy, including TP, FP, FN, TN, and F1-score.
@@ 4.3.2. RQ2: Performance of LLM-SmartAudit Across Different Strategies
Table 3. Performance Comparison of Different Detection Methods on Real-World Dataset
@@
4.3.3. RQ3: Evaluation Using Real-world Dataset
To address RQ3, we present a comprehensive evaluation of LLM-SmartAudit system using the Real-world dataset. The LLM-SmartAudit system design allocates approximately 1,000 tokens for prompt engineering, excluding the context of code snippets. Given GPT-3.5’s default token limit of 4,096, only 3,000 tokens remain available for contract content within the prompt. For the Labeled dataset, the analyzed contracts are less than 3,000 tokens in length. However, with the Real-world datasets, 533 contracts exceed 3,000 tokens in length. Consequently, only GPT-4 was utilized for this evaluation due to its higher token support of up to 128,000 tokens in context. The evaluation focuses on two primary metrics: TPs and recall, with results compared against audit reports [^3] https://github.com/ZhangZhuoSJTU/Web3Bugs/tree/main/reports. Table 3 presents the evaluation results.
@@ 4.3.3. RQ3: Evaluation Using Real-world Dataset
While earlier works such as David et al. (David et al., 2023) and GPTScan (Sun et al., 2024b) demonstrated the potential of GPT-based vulnerability detection tools, our research significantly broadens the scope and improves upon their findings. However, due to the unavailability of aforementioned tools and insufficient information for replication, this study relies on their published statistics for comparison. David et al. reported a recall rate of 39.73%, detecting 58 out of 146 vulnerabilities using a combination of Claude and GPT. GPTScan analyzed 232 vulnerabilities across 72 projects, correctly identifying 40 true positives. In contrast, our study substantially broadens the scope, examining 1,408 vulnerabilities across 102 projects. This expanded scope allows for a more comprehensive assessment of GPT-based tools’ capabilities in real-world scenarios. Our results show a notable improvement, with 672 out of 1,408 vulnerabilities detected, yielding an overall recall rate of 47.73%.
@@
4.3.4. RQ4: Newly Discovered Vulnerabilities
In response to RQ4, our methods successfully identified 11 vulnerabilities across 4 different types that were not detected in the audit reports from real-world datasets. These findings have been submitted to the Code4rena community for verification. Table 4 provides the list of these newly discovered vulnerabilities.
Table 4. Summary of Newly Discovered Vulnerabilities
Vulnerability Affected Locations
Unlimited Token Approval supplyTokenTo() in SushiYieldSource.sol; .approve() in NFTXStakingZap.sol; _par.approve() in PARMinerV2.sol
Insufficient Input Validation setTransferRatio() in sYETIToken.sol
Improper Partial Withdrawals withdraw() in yVault.sol; processWithdraw() in synthVault.sol
Unchecked External Calls earn(), withdraw() in yVault.sol; vault.addValue(), getUnifiedAssets() in IndexTemplate.sol; collectEarnings(), _push(), _pullUniV3Nft() in UniV3Vault.sol
@@ 4.4. Related Work
David et al. (David et al., 2023) examined the efficacy of LLMs, like GPT-4 and Claude, in auditing DeFi smart contract security. Their study employed a binary classification approach, asking the LLMs to determine whether a contract is vulnerable. Although GPT-4 and Claude demonstrated high true positive rates, they also exhibited significant false positive rates. The researchers highlighted the substantial evaluation cost, approximately 2,000 USD, for analyzing 52 DeFi attacks. Chen et al. (Chen et al., 2023b) performed a comparative analysis of GPT’s smart contract vulnerability detection capabilities against established tools. Their results revealed varying GPT effectiveness across common vulnerability types, encompassing 8 types compared to the 10 in our study. Sun et al. (Sun et al., 2024b) evaluated GPT’s function vulnerability matching using a binary response format (’Yes’ or ’No’) for predefined scenarios. They also highlighted potential false positives due to GPT’s inherent limitations. While their study found no significant improvements with GPT-4, our research and others have demonstrated GPT-4’s enhanced detection capabilities.
In addition to SOTA commercial products, open-source alternatives have been considered for smart contract analysis. Shou et al. (Shou et al., 2024) integrates Llama-2 model into the process of fuzzing to detect vulnerabilities in smart contracts, aiming to address inefficiencies in traditional fuzzing methods. However, this approach’s efficacy depends on LLMs’ accurate and nuanced understanding of smart contracts, and it confronts challenges in complexity, cost, and dependence on static analysis. Sun et al. (Sun et al., 2024a) explored open-source tools such as Mixtral and CodeLlama against GPT-4 for detecting smart contract vulnerabilities. They discovered that GPT-4, leveraging its advanced Assistants’ functionalities to effectively utilize enhanced knowledge and structured prompts, significantly outperformed Mixtral and CodeLlama. However, the assessment was limited to demonstrations from the Replicate website, potentially not fully representing these LLMs’ capabilities.
5. Discussions
5.1. Summary of Findings
@@
5.2. Threats of Validity
Our proposed system has the following potential limitations:
@@ 5.2. Threats of Validity
This section has provided a comprehensive overview of the LLM-SmartAudit system, highlighting its strengths and limitations. These insights form our study, offering a nuanced understanding of the potential and challenges of using LLMs in smart contract analysis.
6. Conclusions